OpenStudy (itsbribro):

What is the equation of a line that is parallel to −3x+4y=4 and passes through the point (4,0)?

OpenStudy (3mar):

For parallel lines: "All parallel lines have the same slope!" And for the equation of the line: take that!

OpenStudy (heart_offire):

im sorry, i have no idea how to do this....lol i dont even know where to start

OpenStudy (itsbribro):

sameee!!!

OpenStudy (gunsodo):

More hints, if same slope... -3x+4y=??? if it also passes 4,0

OpenStudy (3mar):

Did not see my comment, @itsbribro ??

OpenStudy (itsbribro):

yes but I don't understand what you mean

OpenStudy (itsbribro):

@Candy13106

OpenStudy (3mar):

Step by step... −3x+4y=4 Can you find out the slope of this line!?

OpenStudy (itsbribro):

yes

OpenStudy (3mar):

What is it?

OpenStudy (itsbribro):

3/4

OpenStudy (3mar):

\[\Huge\color{red}\checkmark\] So you have the slope. Just we need a point to form the equation of our line. "and passes through the point (4,0)" Can you use the formula I sent and plug in your input into it?

OpenStudy (itsbribro):

is it 0=3/4*4+b

OpenStudy (itsbribro):

...

OpenStudy (itsbribro):

i have no clue.. i'm sorry

OpenStudy (3mar):